OpenStudy (anonymous):

if the point (2/3 y) is on the unit circle what is y? a. 2/3 b. 1/3 c. sqrt 5/3 d. 4/9

OpenStudy (kc_kennylau):

All points on the unit circle satisfy the property \(x^2+y^2=1\)

OpenStudy (kc_kennylau):

Solve \(\Large\left(\dfrac23\right)^2+y^2=1\) to find the answer

OpenStudy (anonymous):

idk if its sqrt of 5/3 or 4/9 because y^2 +4/9=1

OpenStudy (kc_kennylau):

Continue please your work

OpenStudy (anonymous):

so its 4/9

OpenStudy (kc_kennylau):

No it's not

OpenStudy (kc_kennylau):

Step by step please?

OpenStudy (kc_kennylau):

\[\Large\begin{array}{rcl} \left(\frac23\right)^2+y^2&=&1\\ \frac49+y^2&=&1\\ y^2&=&\frac59\\ y&=&\frac{\sqrt5}3 \end{array}\]

OpenStudy (kc_kennylau):

How did you get 4/9 as the answer?

OpenStudy (anonymous):

ohhh ok i didnt do the second half